OpenStudy (secret-ninja):

A polynomial function can be written as (x - 1)(x - 4)(x + 7). What are the x-intercepts of the graph of this function? (1, 0), (4, 0), (7, 0) (-1, 0), (-4, 0), (-7, 0) (1, 0), (4, 0), (-7, 0) (-1, 0), (-4, 0), (7, 0)

OpenStudy (secret-ninja):

I think its D. :p

OpenStudy (ccswims):

Find the zeros x - 1 = 0 x - 4 = 0x + 7 = 0

OpenStudy (ccswims):

sorry that got mixed up

OpenStudy (ccswims):

x - 1= 0 x - 4 = 0 x + 7 = 0

OpenStudy (ccswims):

for example, the first one is x - 1 = 0 then do inverse operations, x + 1 = 0 + 1 So one of the x-intercepts are (1, 0)

OpenStudy (ccswims):

If you do it that way for the others you should actually get C as your answer

OpenStudy (secret-ninja):

Ah, so they weren't negative. Thank you! ^-^